# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USAthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USA

### v1

@@ E.3 Analysis of experimental results
Appendix F No-jump backaction in the SWS gate
@@
Appendix G CZ induced coupler heating
Since the transmon coupler mode $c$ is temporarily populated during the application of the CZ gate it is important to ensure that all the population is transferred back to the control dual-rail before the gate ends, failing to do so will dephase the target qubit. In this section we describe the experimental protocol to measure the residual coupler excitation at the end of the CZ gate. We do this by preparing our system in the state $|01g10\rangle$ and repeatedly applying the CZ gate followed by a slow selective $X$ gate on the target qubit. This selective $X$ gate only swaps excitation between cavities $b_{1}$ and $b_{2}$ when the coupler is in the $|g\rangle$ state and implements an identity on the target qubit when the coupler is in the $|e\rangle$ state. This pulse sequence maps the residual coupler excitation onto the target qubit $|01\rangle$ state population. At the end of the sequence we measure the two dual-rail qubits. We process the data in two ways. First, we trace out the control qubit and measure the population of the target qubit in the state $|10\rangle$. This encodes the unpostselected residual coupler excitation at the end of $N$ CZ gate applications. We observe a small but measurable increase in the coupler population of  $0.014\%$ per gate application (see Fig 8 b). Next, we use the same dataset but postselect on the measurement outcomes such that the control qubit is assigned to be in the codespace. In this case we observe no measurable increase of the target qubit $|10\rangle$ population after 64 applications of the gate. Its important to note here that because of the excitation preserving property of the gate whenever we inadvertently leave the coupler in the excited state we will necessarily observe the control qubit in the vacuum sate. This allows us to postselect on this error channel. Lastly, we also compare this measurement to the case when we replace the CZ gate with a delay of equivalent duration and find that we see no measurable difference between the two sequences when the control qubit is postselected to be in the code space.
@@ Appendix G CZ induced coupler heating
Appendix H Correlated leakage and coupler excitation in the SWS gate
@@
Appendix I Measuring bit-flip errors
@@
Appendix J Derivation of gate parameters
@@
One thing to note is that for a given set of hardware parameters, $(g_{ac},\chi_{bc})$, there are multiple sets of pulse parameters, $(t_{\mathrm{SWAP}},t_{\mathrm{wait}},\phi)$ which can enact a $\mathrm{CZ}$ gate. To simplify this, we can choose to fix one of the parameters. In [39], they choose to fix $\phi=\pi$, and calibrate the two pulse timing parameters. We instead choose to fix $t_{\mathrm{SWAP}}=\pi/g_{ac}$, taking $g_{ac}$ to be real and positive. (Experimentally, this means we only need to calibrate a delay time and a phase, the easiest parameters to sweep).
@@
$t_{\mathrm{wait}}$ $=\frac{\pi}{|\chi_{bc}|}-\frac{\pi}{g_{ac}}$ (33)
@@
$+2\arctan{\left(-\frac{\sqrt{g_{ac}^{2}+\chi_{bc}^{2}}}{\chi_{bc}}\cot{\left(\frac{\pi\sqrt{g_{ac}^{2}+\chi_{bc}^{2}}}{2g_{ac}}\right)}\right)}$ ) (35)
With these parameters, the gate can be modeled through three different piece-wise Hamiltonians:
@@
J.1 Deriving $t_{\mathrm{wait}}$
@@
J.2 Deriving $\phi_{\mathrm{SWAP}}$
As may be expected from the form of the analytical expression for $\phi_{SWAP}$ in Eq. 33, this derivation will require some geometry and trigonometry. It is helpful to refer to Fig. 10. The angle $\phi_{3}$ is related to $\phi_{\mathrm{SWAP}}$ as $\phi_{3}=\pi-\phi_{\mathrm{SWAP}}$. We also know the value of $t_{\mathrm{wait}}$ and so $\phi_{2}=\chi_{bc}t_{\mathrm{wait}}$. If we can calculate $\phi_{2}$ then we can use the fact $\phi_{1}+2\phi_{2}+\phi_{3}=2\pi$ to calculate $\phi_{3}$ and thus $\phi_{\mathrm{SWAP}}$. We start by noting that the evolution during $\mathcal{H}_{1}$ (and $\mathcal{H}_{3}$) is an ellipse in the XY plane. On the Bloch sphere, these amount to detuned Rabi oscillations but when projected onto the XY plane these become elliptical. We define the major axis of this ellipse to be $2y_{0}$ and the minor axis to be $2x_{0}$. It can be shown that the ratio $y_{0}/x_{0}=\chi_{bc}/\Omega$.
@@
Appendix K QuTiP Master equation simulations
We simulate expected gate performance by simulating evolution under the three piecewise Hamiltonians described in 36 (i.e. with the square pulse simplification), modelling all 5 modes $a_{1},a_{2},c,b_{1},b_{2}$ and their respective collapse operators, modelling white-noise dephasing and excitation loss. The loss rates for these collapse operators are obtained from our measured hardware coherence times.
These simulations capture the effects of no-jump back-action, expected Z-frame updates, correlated errors, and verify that our expressions for derived gate parameters in 33 are exact.
@@
Figure 11: Evolution of the dual-rail basis states |$\ket{0_{L}1_{L}}$ (left) and |$\ket{0_{L}0_{L}}$ (right) during a the SWS gate, obtained from our QuTiP simulations.
Appendix L Fundamental limits to SWS gate performance
@@
SWS error rate Error scaling
$p_{e}^{\mathrm{control}}$ $\frac{1}{G}\frac{1}{\alpha_{c}T_{c}^{1}}$
$p_{e}^{\mathrm{target}}$ $\frac{1}{\alpha_{c}T_{c}^{1}}$
$p_{Z}^{\mathrm{control}}$ $\frac{1}{G}\frac{1}{\alpha_{c}T_{c}^{\phi}}$
$p_{Z}^{\mathrm{target}}$ $G\times\frac{1}{\alpha_{c}T_{c}^{\phi}}$

### v2

@@ F.3 Analysis of experimental results
Appendix G No-jump backaction in the SWS gate
@@
Appendix H CZ induced coupler heating
Since the transmon coupler mode $c$ is temporarily populated during the application of the CZ gate it is important to ensure that all the population is transferred back to the control dual-rail before the gate ends, failing to do so will dephase the target qubit. In this section we describe the experimental protocol to measure the residual coupler excitation at the end of the CZ gate. We do this by preparing our system in the state $|01g10\rangle$ and repeatedly applying the CZ gate followed by a slow selective $X$ gate on the target qubit. This selective $X$ gate only swaps excitation between cavities $b_{1}$ and $b_{2}$ when the coupler is in the $|g\rangle$ state and implements an identity on the target qubit when the coupler is in the $|e\rangle$ state. This pulse sequence maps the residual coupler excitation onto the target qubit $|10\rangle$ state population. At the end of the sequence we measure the two dual-rail qubits. We process the data in two ways. First, we trace out the control qubit and measure the population of the target qubit in the state $|10\rangle$. This encodes the unpostselected residual coupler excitation at the end of $N$ CZ gate applications. We observe a small but measurable increase in the coupler population of  $0.014\%$ per gate application (see Fig 9 b). Next, we use the same dataset but postselect on the measurement outcomes such that the control qubit is assigned to be in the codespace. In this case we observe no measurable increase of the target qubit $|10\rangle$ population after 64 applications of the gate. Its important to note here that because of the excitation preserving property of the gate whenever we inadvertently leave the coupler in the excited state we will necessarily observe the control qubit in the vacuum sate. This allows us to postselect on this error channel. Lastly, we also compare this measurement to the case when we replace the CZ gate with a delay of equivalent duration and find that we see no measurable difference between the two sequences when the control qubit is postselected to be in the code space.
@@ Appendix H CZ induced coupler heating
Appendix I Correlated leakage and coupler excitation in the SWS gate
@@
Appendix J Measuring bit-flip errors
@@
Appendix K Derivation of gate parameters
@@
One thing to note is that for a given set of hardware parameters, $(g_{ac},\chi_{bc})$, there are multiple sets of pulse parameters, $(t_{\mathrm{SWAP}},t_{\mathrm{wait}},\phi)$ which can enact a $\mathrm{CZ}$ gate. To simplify this, we can choose to fix one of the parameters. In [40], Rosenblum et al. choose to fix $\phi=\pi$, and calibrate the two pulse timing parameters. We instead choose to fix $t_{\mathrm{SWAP}}=\pi/g_{ac}$, taking $g_{ac}$ to be real and positive. (Experimentally, this means we only need to calibrate a delay time and a phase, the easiest parameters to sweep).
@@
$t_{\mathrm{wait}}$ $=\frac{\pi}{|\chi_{bc}|}-\frac{\pi}{g_{ac}},$ (42)
@@
$\quad+2\arctan{\left(-\frac{\sqrt{g_{ac}^{2}+\chi_{bc}^{2}}}{\chi_{bc}}\cot{\left(\frac{\pi\sqrt{g_{ac}^{2}+\chi_{bc}^{2}}}{2g_{ac}}\right)}\right)},$ (44)
derived below. With these parameters, the gate can be modeled through three different piece-wise Hamiltonians:
@@
K.1 Deriving $t_{\mathrm{wait}}$
@@
K.2 Deriving $\phi_{\mathrm{SWAP}}$
As may be expected from the form of the analytical expression for $\phi_{\mathrm{SWAP}}$ in Eq. 42, this derivation will require some geometry and trigonometry. It is helpful to refer to Fig. 11. The angle $\phi_{3}$ is related to $\phi_{\mathrm{SWAP}}$ as $\phi_{3}=\pi-\phi_{\mathrm{SWAP}}$. We also know the value of $t_{\mathrm{wait}}$ and so $\phi_{2}=\chi_{bc}t_{\mathrm{wait}}$. If we can calculate $\phi_{2}$ then we can use the fact $\phi_{1}+2\phi_{2}+\phi_{3}=2\pi$ to calculate $\phi_{3}$ and thus $\phi_{\mathrm{SWAP}}$. We start by noting that the evolution during $\mathcal{H}_{1}$ (and $\mathcal{H}_{3}$) is an ellipse in the XY plane. On the Bloch sphere, these amount to detuned Rabi oscillations but when projected onto the XY plane these become elliptical. We define the major axis of this ellipse to be $2y_{0}$ and the minor axis to be $2x_{0}$. It can be shown that the ratio $y_{0}/x_{0}=\chi_{bc}/\Omega$.
@@
Appendix L QuTiP Master equation simulations and CZ gate error budget
Figure 12: Evolution of the dual-rail basis states |$\ket{0_{L}1_{L}}$ (left) and |$\ket{0_{L}0_{L}}$ (right) during the SWS gate, obtained from our QuTiP simulations.
Error channel Error state(s) after gate Probability (QuTiP) Probability (at short gate depth)
Photon loss on control DR |$\ket{00g01},\ket{00g10}$ 0.337% 0.400(4)%
Photon loss on target DR |$\ket{10g00},\ket{01g00}$ 0.0982% 0.096(4)%
Photon loss on both DRs |$\ket{00g00}$ 0.0003% -
Control DR excitation stuck in coupler mode |$\ket{00e10},\ket{00e01}$ 0.006% -
Only one excitation remains in coupler mode |$\ket{00e00}$ 0.0014% -
Control DR dephasing $Z_{c}\ket{\psi_{\mathrm{ideal}}}$ 0.0169% 0.039(1)%
Target DR dephasing $Z_{t}\ket{\psi_{\mathrm{ideal}}}$ 0.0048% 0.0112(9)%
Control and target DR dephasing $Z_{c}Z_{t}\ket{\psi_{\mathrm{ideal}}}$ 0.0001% -
No error |$\ket{\psi_{\mathrm{ideal}}}$ 99.5351% -
Table 5: Estimated error budget for a single CZ gate, based on QuTiP simulations and measured hardware coherence times. The rightmost column compares to experimentally measured values at short numbers of gate repeates.
We simulate expected gate performance by simulating evolution under the three piecewise Hamiltonians described in 45 (i.e. with the square pulse simplification), modelling all 5 modes $a_{1},a_{2},c,b_{1},b_{2}$ and their respective collapse operators, modelling white-noise dephasing and excitation loss. The loss rates for these collapse operators are obtained from our measured hardware coherence times.
These simulations capture the effects of no-jump back-action, expected Z-frame updates, correlated errors, and verify that our expressions for derived gate parameters in 42 are exact.
@@
We can use this QuTiP simulation to numerically simulate the expected error budget for our gate, given measured hardware coherence times, and compare it against some of our measured gate errors. For this, we use the coherence times listed in Tab. 2 and 3, taking the longer echo times as our dephasing times. The simulated error budget is shown in Tab. 5 and allows us to estimate the probabilities for different CZ gate error channels that are too small to measure directly in experiment.
Appendix M Fundamental limits to SWS gate performance
SWS error rate Error scaling
$p_{e}^{\mathrm{control}}$ $\frac{1}{G}\frac{1}{\alpha_{c}T_{c}^{1}}$
$p_{e}^{\mathrm{target}}$ $\frac{1}{\alpha_{c}T_{c}^{1}}$
$p_{Z}^{\mathrm{control}}$ $\frac{1}{G}\frac{1}{\alpha_{c}T_{c}^{\phi}}$
$p_{Z}^{\mathrm{target}}$ $G\times\frac{1}{\alpha_{c}T_{c}^{\phi}}$